# Large Language Models:

### v1

@@ 5.3.1 Linear Regression with LLMs
Second, she may regress the linked variable on the economic concept
@@
The bias of the plug-in regression coefficients $\beta$ and $\alpha$ depends on how the LLM’s measurement error $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ varies across text pieces. These are well-known results, dating back to BoundEtAl(94)-ValidationData, that we restate in our framework.
@@
Consider the research context $Q(\cdot)$ and assume that $q_{r}^{T}(t_{r})=q_{r}^{T\mid D}(t_{r})$ for all $r\in\mathcal{R}$.
  1. 1.
@@
In other words, when the economic concept is the dependent variable, the bias of the plug-in regression coefficient $\beta$ is summarized by the extent to which the LLM’s measurement error $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ covaries with the linked variable $W_{r}$ in the research context. Measurement error in the LLM’s label induces an omitted variable bias in the plug-in regression coefficient. While its expression is more complicated when the economic concept is the covariate, the bias of the plug-in regression coefficient $\alpha$ can again be summarized by how the economic concept covaries with the LLM’s label.
To concretely illustrate why researchers must be appropriately skeptical of claims that the labels of LLMs are as-good-as ground truth in economic settings, we return to our earlier examples of financial news headlines and congressional legislation. If they are equivalent to ground truth measurements, it must be that the parameter estimates based on LLM labels should not be sensitive to specific implementation choices, such as the particular choice of large language model or the specific prompt engineering strategy. By contrast, Proposition 3 implies that variation in the estimates across models and prompting strategies must be driven by variation in the associated measurement error. If not—if the parameter estimates are sensitive to these LLM implementation choices—then clearly the labels of some of these LLMs cannot be treated as if they are ground truth.
5.3.2 Assessing Measurement Error in Financial News Headlines
@@
We return to the dataset from Section 4.3 of financial news headlines for $6,000$ publicly traded stocks from 2009-2020. We now focus on all financial news headlines published in 2019. We observe the text of each headline $r$, its publication date, and the ticker of the stock it refers to. Onto each financial news headline $r$, we merge the stock’s realized returns within various windows after the publication date $W_{r}$ (e.g., 1 day, 5 days, and 10 days after publication) as well as its lagged returns before the publication date (e.g., 1 day, 2 days and 3 days before publication) (WRDSBetaSuite).
@@
Financial news headlines express various economic concepts that we could measure and relate to realized stock returns. In this exercise, we focus on one simple economic concept: is the headline positive, negative or neutral news for the company? As an example, consider the headline
Bank of New York Mellon Q1 EPS \$0.94 Misses \$0.97 Estimate, sales \$3.9B Miss \$4.01B Estimate.
Surely this is negative news for Bank of New York Mellon since it indicates the bank underperformed relative to market expectations. Now consider the headline
Disney And Charter Communications Announce Comprehensive Distribution Agreement; New Multi-Year Deal Delivers Full Suite Of Sports, News And Entertainment Networks From Walt Disney Television And ESPN To Spectrum Customers.
This is reasonably labeled as positive news for Charter Communications since the distribution agreement with Disney enables Charter Communications to expand its offerings to existing and potential customers. While we could in principle read every single financial news headline published in 2019, this would obviously be a painstaking process.
@@
So it is natural to instead turn to LLMs to solve this text processing problem. For our own analysis, we take each financial news headline $r$ and prompt LLMs to label each news headline as positive, negative or neutral news for the company it refers to, as well as score the magnitude of this assessment and the LLM’s confidence in the label, $\widehat{V}_{r}=\widehat{m}(r;t)$. This requires making several practical choices (as does any effort to solve any text processing problem with an LLM): what specific LLM to use? What prompt engineering strategy? Economists happy to treat LLM labels as ground truth should have their faith shaken if alternative answers to these practical questions lead to meaningfully different labels and downstream parameter estimates—this would indicate non-ignorable measurement error for this economic concept.
@@
To examine this possible sensitivity, we separately prompt GPT-3.5-Turbo, GPT-4o, and GPT-4o-mini to label each financial news headline $r$, using nine alternative prompts to each model. Our two base prompts provide the LLM with the text of the headline $r$ and ask it to label whether this news is positive, negative, or neutral for the company $V_{r}$; we also ask the LLM to provide its confidence and a magnitude in the label. Our two base prompts differ in how they ask the model to format its reply: filling-in-the-blanks text or a structured JSON (JavaScript Object Notation) object. The exact wordings of our base prompts are provided in Appendix Figure 13b.
@@
To round out our list of nine different prompts, we also vary them in two other ways as well. These variations are motivated by existing computer science research on prompting LLMs, which provides a large menu of prompt engineering strategies that can be appended to a base prompt and have been found to meaningfully affect the performance of LLMs on natural language processing tasks (LiuEtAl(23); wei2023chainofthoughtpromptingelicitsreasoning; white2023promptpatterncatalogenhance; chen2024unleashingpotentialpromptengineering). First, we append to the prompt a request to the LLM to adopt one of four different “personas,” like “knowledgeable economic agent” or “expert in finance” (see Appendix Figure 14d for the exact wording). We also append three prompt modifiers that ask the LLM to “think carefully” or “think step-by-step” and provide an explanation for its answer (see Appendix Figure 15c for the exact wording). Altogether, for each financial news headline $r$, we obtain labels $\widehat{V}_{r}^{m,p}$ associated with three different LLMs $m$ and nine different prompting strategies $p$.
We first calculate the pairwise agreement in the labels produced by alternative prompting strategies $p$. The results, summarized in Figure 3c separately for each model $m$, show substantial variation in the pairwise agreement in the labels produced. For GPT-3.5-Turbo, our base prompt with fill-in-the-blank formatting only produces the same label as our base prompt with JSON formatting on $67\%$ of financial news headlines. Commanding GPT-3.5-Turbo to behave as a “knowledgeable economic agent” produces the same label as commanding it to behave as a “successful trader” on $91.6\%$ of financial news headlines. We see similar variation across pairs of prompting strategies for GPT-4o and GPT-4o-mini. Prompting GPT-4o to “think step-by-step” produces the same label as our base prompt with fill-in-the-blank formatting on $89.2\%$ of headlines; whereas prompting GPT-4o to answer like an “expert in finance” versus an “expert in the economy” leads to agreement on $96.3\%$ of headlines. Across models, there appear to be no consistent patterns in which pairs of prompting strategies tend to have the most agreement.
Of course, economists don’t care about prompt responses per se; we care about what happens when we use these prompt responses to estimate economically relevant parameters. To what extent does this variation in LLM output matter for downstream parameter estimation?
@@
To investigate this, for each model $m$ and prompt $p$, we next regress the realized returns of each stock within 1-day, 5-days and 10-days after the headline’s publication date $Y_{r}$ on binary indicators for whether the LLM’s labels $\widehat{V}_{r}^{m,p}$ are positive versus negative (with neutral the omitted reference category), controlling for the model’s reported magnitudes and lagged realized returns. We report separately the coefficients $\widehat{\beta}_{m,p}$ on whether the LLM labels the headline as positive and whether the LLM labels it as negative, as well as the coefficient’s associated t-statistics with standard errors clustered at the date and company level.
Figure 4 summarizes the variation in the resulting t-statistics across models and prompts in the regression of the LLM’s labels $\widehat{V}_{r}^{m,p}$ against 1-day, 5-day and 10-day realized returns. Table 2(b) summarizes the variation in the coefficient estimates across models and prompts for all horizons of realized returns. In Appendix Figures A3 and Appendix Table 3(b), we show results that controls for the LLM’s assessed confidence; results are qualitatively similar.
For the exact same set of financial news headlines, simply adjusting the prompting strategy or using a different LLM can yield remarkably different results. For example, looking at the coefficient on the positive headline indicator and 1-day realized returns, the results range from a t-statistic of nearly $-6$ up to $+2$ (Figure 4). Furthermore, the variation in coefficients is meaningful in terms of economic magnitudes even when the estimates share the same sign. This problem is even worse for returns on longer horizons, as the bottom panels of Figure 4 show. In other words, there are many combinations of prompt and model that produce entirely different directions and magnitudes of the relationship between the positive/negative label and realized returns.
5.3.3 Assessing Measurement Error in Congressional Legislation
@@
We next return to the Congressional legislation data described earlier. For each bill, we observe the text of its description $r$ as well as a collection of additional linked economic variables $W_{r}$, such as the party affiliation of the bill’s sponsor, whether the bill originated in the Senate, and an ideological score – the DW1 roll call voting record – of the bill’s sponsor. A natural political-economy question might be how these variables shape the topic of each bill, $V_{r}$. Could we draw valid inferences using LLMs to collect those labels rather than relying on human annotation?
@@
For our analysis, we randomly select $10,000$ Congressional bills and separately prompt GPT-3.5-Turbo and GPT-4o to label each Congressional bill for its policy area using alternative prompting strategies, including base prompts that modify the requested format, persona modifications, chain-of-thought modifications, and even few-shot examples (see Appendix E.4 for the specific prompts we used). Consequently, for each Congressional bill description $r$, we obtain labels $\widehat{V}_{r}^{m,p}$ associated with two different large language models $m$ and twelve prompting strategies $p$. Appendix Figure A4 calculates the pairwise agreement in the labels produced by alternative prompting strategies $p$. We again find substantial variation in the pairwise agreement in the labels produced, and there appear to be no consistent patterns in which pairs of prompting strategies tend to have the most agreement.
@@
We then separately regress the labeled economic concept — in this case, the policy topic area of the bill — against linked covariates $W_{r}$, separately reporting the coefficients $\widehat{\beta}_{m,p}$ and their associated t-statistics. Figure 5 summarizes the variation in the resulting t-statistics across models and prompts: each column shows the result of using LLM labels $\widehat{V}_{r}^{m,p}$ for alternative bill topics as the dependent variable (i.e., Health, Banking, Finance and Domestic Commerce, Defense, Government Operations, and Public Lands and Water Management), and each row shows the result using a different linked covariate $W_{r}$ (i.e., whether the bill’s sponsor was a Democrat, whether the bill originated in the Senate, and the DW1 score of the bill’s sponsor). Table 3 summarizes the coefficient estimates across models and prompts for each choice of labeled policy topic and the covariate. As was the case with financial news headlines, for each combination of labeled policy topic and linked covariate in the Congressional bills dataset we see substantial variability in the resulting t-statistics across different LLMs and prompting strategies.
5.4 Recommendations for Empirical Practice
When economists see an LLM ace the SAT or GRE we must resist the natural urge to anthropomorphize the algorithm and assume it is capable of performing all sorts of other tasks that a human who performs well on standardized tests could be expected to carry out. We cannot simply plug the LLM’s output into an economics research pipeline and “set it and forget it,” assuming the implicit claims that LLMs are providing ground-truth output. Viral examples of the capabilities of LLMs and benchmark evaluations do not justify the strong guarantee needed for estimation problems.
The good news is that there is a solution. The key to safely using this new technology is to actually look back to old lessons in econometrics. Articulating statistical assumptions about the structure of measurement error in LLMs appears challenging — after all, these are complicated algorithms trained on unknown datasets. Yet a classic solution can still be applied: invest in collecting the gold standard label $V_{r}$ on a small validation sample and use it to de-bias the plug-in estimate based on the LLM labels $\widehat{V}_{r}$. The virtues of validation data in dealing with measurement error is well-known – see, for example, classic work such as BoundKrueger(91); BoundEtAl(94)-ValidationData; BoundEtAl(01)-HandbookChapter and LeeSepanski(95)-ErrorsInVariable; chen2005measurement. It has been recently revived in machine learning; see WangMcCormickLeek(20); AngelopoulosEtAl(23)-PPI; EgamiEtAl(24)-ImperfectSurrogates. The corollary of this classic solution is that with any application in which the researcher cannot collect validation data, the researcher really should not be using the LLM’s labels for estimation problems.
In the rest of this section, we demonstrate the value of debiasing LLM labels in the context of linear regression. We review the mechanics of debiasing a linear regression in which the LLM is the dependent variable and provide intuition about when it will work well based on asymptotic arguments. We then show this empirically in finite samples using our earlier example on Congressional legislation before turning to briefly discuss how these results carry over to the case of using the LLM’s output as an explanatory rather than dependent variable.
5.4.1 De-Biasing LLM Labels: Conceptual Results for Linear Regression
@@
Let us return to the case where the researcher wishes to regress the economic concept $V_{r}$ on the linked variables $W_{r}$, but relies on the LLM labels instead and reports the results of estimating the plug-in regression $\widehat{m}(r;t)=W_{r}^{\prime}\beta+\widetilde{\epsilon}_{r}$. As noted above, the bias of the plug-in regression coefficient $\beta$ is summarized by the extent to which the LLM measurement error $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ covaries with the linked variable $W_{r}$ in the research context. Our discussion here extends regressing the linked variable on the economic concept (Equation 11); see Appendix C.2 for details.
Proposition 3 described above has two implications that both hinge on collecting the gold standard label $V_{r}$ on a small validation sample. More specifically, on a random subset of the researcher’s dataset, we will now assume that the researcher collects the label $V_{r}$. We refer to the collection of text pieces on which the researcher now observes $(r,W_{r},\widehat{m}(r;t),V_{r})$ as the validation sample, and we refer to the remaining text pieces on which she observes $(r,\widehat{m}(r;t),V_{r})$ as the primary sample. Indeed, in many empirical applications that use LLMs to solve an estimation problem, it is already common for researchers to collect such a validation sample (e.g., durvasula2024distillingdatalargelanguage; HansenKazinnik(24)). Often the validation data is used for purposes of reporting the overall accuracy of the LLM’s labels for the economic concept; although, as we have shown, this is not sufficient to ensure valid inference for whatever downstream parameters the labels are used to estimate. The real (often unrealized) value of the validation sample comes from enabling the researcher to empirically estimate the bias associated with the plug-in regression; that is, the researcher can estimate $\widehat{\lambda}_{\Delta\mid W}$ by forming $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ on the validation sample and regressing it on the linked variable $W_{r}$. By making the bias of the LLM estimable, there are two implications for empirical researchers.
@@
First, the validation sample provides the researcher with a direct target to optimize when fine-tuning or prompt engineering a given LLM. Since her goal is to ultimately run some downstream regression, her goal is to obtain quality downstream estimates rather than simply maximizing the accuracy of the resulting labels — which, as we saw in the Congressional bills application, can provide a misleading picture. For any choice of LLM $m$ and prompt $p$, the researcher can estimate the bias of the plug-in regression $\widehat{\lambda}_{\Delta\mid W}^{m,p}$ associated with the labels $\widehat{V}_{r}^{m,p}$; and thereby select the combination that results in the smallest bias.
@@
Second, more importantly, for any choice of an LLM and prompting strategy, the validation sample can be used to bias-correct the plug-in regression constructed on the primary sample. In other words, rather than directly reporting the plug-in regression $\widehat{\beta}$, the researcher can instead report the bias-corrected estimator
@@
This is exceedingly simple to implement: it only involves running two linear regressions, regressing $\widehat{m}(r;t)$ on $W_{r}$ in the primary sample and $\Delta_{r}$ on $W_{r}$ in the validation sample.
As we show in more detail in Appendix C.2, the bias-corrected estimator has desirable theoretical properties. Consider the case in which the researcher’s dataset is a random sample of economically relevant text pieces, a fraction $\rho_{p}$ of all text pieces enter into the primary sample, and a fraction $\rho_{v}$ of all text pieces enter into the validation sample. As the number of economically relevant text pieces grows large, the bias-corrected estimator $\widehat{\beta}^{debiased}$ is consistent for the target regression coefficient $\beta^{*}$. Moreover, it can be shown that $\widehat{\beta}^{debiased}$ is asymptotically normal with limiting variance given by
@@
where $\sigma_{W}$ is the standard deviation of the linked variable $W_{r}$ across all text pieces, $\sigma_{\widehat{V}W}$ is the standard deviation of the product $\widehat{V}_{r}\times W_{r}$, and $\sigma_{\Delta W}$ is defined analogously. Consequently, the precision of the bias-corrected estimator depends on the relative size of the validation sample versus the primary sample as well as the variability of the LLM’s label $\widehat{V}_{r}$ and measurement error $\Delta_{r}$ across text pieces.
Of course, if the researcher invests effort to collect ground-truth labels on a validation sample, a natural question arises: why bother with the possibly mis-measured LLM labels on the primary sample? The researcher could just directly estimate the target regression on the validation sample only and report $\widehat{\beta}^{*}$. To answer that question we can analyze the limiting behavior of the validation-sample only regression estimator: it is also asymptotically normal with limiting variance given by $\sigma^{-4}_{W}\frac{1-\rho_{v}}{\rho_{v}}\sigma^{2}_{VW}$. Comparing this expression with that for the bias-corrected regression, the latter is more precise if
@@
Unsurprisingly this comparison depends on the relative size of the validation sample versus the primary sample. But notice that it further depends on the relative standard deviation of the gold standard label $V_{r}$ and the measurement error $\Delta_{r}$. Importantly, Equation (14) implies that the bias-corrected regression coefficient can be more precisely estimated than the validation-sample-only regression coefficient if the LLM’s labels are sufficiently accurate. [^11] This phenomenon has been documented elsewhere in recent machine learning research, such as (AngelopoulosEtAl(23)-PPI) and associated work, that combines validation data with the outputs of machine learning models to estimate downstream parameters. In this case, the researcher can obtain a possible free-lunch: use the LLM to solve the text processing problem and still deliver precise estimates of the target regression coefficient. Importantly, the LLM outputs are not substitutes for gold standard labels; instead, they can serve to cheaply amplify the validation data.
Of course whether this occurs in practice will depend on the specific empirical application. We turn next to illustrating the performance of the bias-corrected regression coefficient in finite samples for our Congressional legislation example.
5.4.2 Monte Carlo Simulations based on Congressional Legislation
@@
The Congressional legislation data described earlier is particularly well-suited to illustrate the performance of the bias-corrected regression because we observe ground-truth labels $V_{r}$ of each bill’s policy topic. The Congressional Bills Project trained teams of human annotators to label the description of each Congressional bill $r$ for its major policy topic area $V_{r}$, describing whether the bill falls into one of twenty possible policy areas such as health or defense. This human labeling process was quite labor intensive. The Congressional Bills Project states that all annotators were trained for a full academic quarter before beginning this task (CAPCodebook).
We start off by reporting a surprising twist on the results reported in the previous section, where we found that plugging in LLM labels for each bill’s policy topic as the dependent variable in a regression led to substantial biases and variation across possible choices of LLM and prompting strategy. The surprising twist is that we observe this substantial variability in downstream parameter estimates even though each LLM and prompt leads to a label that is nearly equally accurate for the ground-truth label $V_{r}$ on average.
Figure 6 plots the accuracy of the LLM label $\widehat{V}_{r}^{m,p}$ for the ground-truth label $V_{r}$ across all combinations of LLMs $m$ and prompting strategies $p$. While GPT-4o tends to be a bit more accurate than GPT-3.5-Turbo, for a given LLM, accuracy is remarkably invariant to prompting strategy. Yet we have seen that for a given LLM the downstream parameter estimates in a regression using these labels varies enormously across different prompting strategies.
To what extent can collecting a small amount of validation data address this problem? To answer that question we carry out a Monte Carlo simulation exercise to examine how well the plug-in regression and bias-corrected regression recover the target coefficient $\beta^{*}$.
@@
More specifically, for a given policy topic $V_{r}$ (e.g., health, defense, etc.), linked covariate $W_{r}$ (e.g., whether the bill’s sponsor was a Democrat, etc.) and pair of large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ bills from our dataset of $10,000$ bills. On this random sample of $5,000$ bills, we first calculate the plug-in coefficient $\widehat{\beta}$ from regressing $\widehat{V}_{r}^{m,p}$ on the linked variable $W_{r}$. We next randomly reveal the ground-truth label $V_{r}$ on 10% of our random sample of $5,000$ bills, which produces a validation sample. We then calculate the bias-corrected coefficient $\widehat{\beta}^{debiased}$. We repeat these steps for $1,000$ randomly sampled datasets. Across each simulation, we calculate the average bias of the plug-in coefficient and the bias-corrected coefficient for the target regression $\beta^{*}$ associated with regressing the ground-truth label $V_{r}$ on the linked variable $W_{r}$ on all $10,000$ bills. We repeat this exercise for each possible combination of bill topic $V_{r}$, linked covariate $W_{r}$, LLM $m$ (either GPT-3.5-Turbo or GPT-4o) and prompting strategy $p$. This allows us to summarize how the plug-in regression performs against the bias-corrected regression across a wide variety of possible regression specifications and choices of LLM and prompting strategy.

### v3

@@ 4.2.1 Some Evidence of Measurement Error
or the independent variable:
@@
The bias depends on how the LLM’s error $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ varies across text pieces. These are well-known results, dating back to BoundEtAl(94)-ValidationData.
@@
Consider the research context $Q(\cdot)$ and assume $q_{r}^{T}(t_{r})=q_{r}^{T\mid D}(t_{r})$ for all $r\in\mathcal{R}$.
  1. (i)
@@
When the economic concept is the dependent variable, the bias equals the best linear predictor of the LLM’s errors given the covariates. When the economic concept is the independent variable, the bias has a more complex form involving attenuation and correlation between the error and the residual. Knowing the LLM is “accurate” provides no guarantee about these biases. What matters is whether errors correlate with economic variables in the specific regression.
Assessing Measurement Error in Financial News Headlines
@@
We return to the financial news headlines dataset, focusing on headlines published in 2019 for $6,000$ publicly traded stocks. We observe each headline’s text $r$, publication date and stock ticker. We merge realized returns at various horizons (1, 5, and 10 days) after publication and lagged returns before publication (WRDSBetaSuite).
@@
Financial news headlines express various economic concepts that we could measure and relate to realized stock returns. We focus on one: is the headline positive, negative or neutral news for the company? While we could read every single financial news headline published in 2019, this process would be painstaking. It is natural to use LLMs to solve this text processing problem.
@@
We take each financial news headline $r$ and prompt LLMs to label each news headline as positive, negative or neutral news for the company it refers to, $\widehat{V}_{r}:=\widehat{m}(r;t)$. This requires making several practical choices: what specific LLM to use? What prompt engineering strategy? If alternative choices lead to different labels and downstream estimates, this would indicate non-ignorable errors in the LLM outputs.
@@
We prompt GPT-3.5-Turbo, GPT-4o, GPT-4o-mini, GPT-5-mini, and GPT-5-nano to label each financial news headline $r$, using nine alternative prompts to each model. Our two base prompts provide the LLM with the text of the headline $r$ and ask it to label whether this news is positive, negative, or neutral for the company; we also ask the LLM to provide its confidence and a magnitude in the label. Our two base prompts differ in how they ask the model to format its reply: filling-in-the-blanks text or a structured JavaScript Object Notation (JSON) object. The prompts are provided in Appendix E.
@@
We further vary these base prompts in two ways, motivated by popular prompt engineering strategies (LiuEtAl(23); wei2023chainofthoughtpromptingelicitsreasoning; white2023promptpatterncatalogenhance; chen2024unleashingpotentialpromptengineering). First, we request the LLM adopt one of four different “personas,” like “knowledgeable economic agent” or “expert in finance.” Second, we append three prompt modifiers that ask the LLM to “think carefully” or “think step-by-step” and provide an explanation for its answer. For each financial news headline $r$, we obtain labels $\widehat{V}_{r}^{m,p}$ associated with a collection of different LLMs $m$ and prompting strategies $p$.
@@
For each model $m$ and prompt $p$, we regress the realized returns of each stock within 1-day, 5-days and 10-days after the headline’s publication date $Y_{r}$ on the LLM’s labels $\widehat{V}_{r}^{m,p}$, controlling for the model’s reported magnitudes and lagged realized returns. We report the coefficients $\widehat{\beta}_{m,p}$ on whether the LLM labels the headline as positive and whether the LLM labels it as negative, as well as their associated t-statistics with standard errors clustered at the date and company level. Figure 1 and Table 1 summarize the results. Simply changing the prompt or model yields markedly different estimates: many prompt-model combinations that produce different directions and magnitudes of the relationship between the positive/negative label and realized returns. [^8] Appendix Figure A3, shows substantial variation in the pairwise agreement in the labels produced. There appear to be no consistent patterns in which pairs of prompting strategies tend to have the most agreement.
Assessing Measurement Error in Congressional Legislation
@@
We next return to the Congressional legislation data. For each bill, we observe the text of its description $r$ as well as a collection of economic variables $W_{r}$, such as the party affiliation of the bill’s sponsor, whether the bill originated in the Senate, and an ideological score – the DW1 roll call voting record – of the bill’s sponsor. A researcher might reasonably study how these variables shape the topic of each bill, $V_{r}$. Could we use an LLM to collect those labels?
@@
We randomly select $10,000$ Congressional bills and separately prompt GPT-3.5-Turbo, GPT-4o, GPT-5-mini, and GPT-5-nano to label each Congressional bill for its policy area using alternative prompting strategies, including base prompts that modify the requested format, persona modifications, chain-of-thought modifications, and even few-shot examples (see Appendix E for the specific prompts we used).
@@
We regress the labeled economic concept $\widehat{V}_{r}^{m,p}$ — in this case, the policy topic of the bill — against linked covariates $W_{r}$, separately reporting the coefficients $\widehat{\beta}_{m,p}$ and their associated t-statistics. Figure 2 and Table 2 summarize the variation in the resulting estimates across models and prompts. For each combination of labeled policy topic and linked covariate in the Congressional bills dataset, we see substantial variability across different LLMs and prompting strategies. [^9] Appendix Figure A4 calculates the pairwise agreement in the labels produced by alternative prompting strategies $p$. We again find substantial variation in the pairwise agreement in the labels produced. Once again, simply changing the prompt or model yields remarkably different downstream estimates.
4.3 Practical Guidance with Validation Data
Given the evidence on LLM brittleness in Section 2.3, it is difficult to defend the assumption of no measurement error in LLM outputs for estimation problems. The solution is to collect measurements $V_{r}=f^{*}(r)$ on a small validation sample and use them to de-bias the plug-in estimate based on the LLM labels $\widehat{V}_{r}$. The virtues of validation data have been understood for decades (bound1991extent; BoundEtAl(94)-ValidationData; LeeSepanski(95)-ErrorsInVariable). It has been recently revived in machine learning to handle the types of mismeasured covariates and outcomes produced by modern ML models; see for example WangMcCormickLeek(20), AngelopoulosEtAl(23)-PPI, EgamiEtAl(24)-ImperfectSurrogates and carlson2025unifyingframeworkrobustefficient among others.
We illustrate the value of validation data using the familiar case of linear regression. We review the mechanics of debiasing when the LLM label appears as the dependent variable, provide intuition about when it will work well using asymptotic arguments, and demonstrate finite sample performance using Congressional legislation. Our discussion is pedagogical; we refer the readers to the works above for more general settings.
An Illustration with Linear Regression
@@ 4.3 Practical Guidance with Validation Data
Return to the case where the researcher wishes to regress the economic concept $V_{r}$ on the linked variables $W_{r}$, but relies on the LLM labels instead and reports the plug-in regression $\widehat{V}_{r}=W_{r}^{\prime}\beta+\widetilde{\epsilon}_{r}$. Appendix C.2 considers when the economic concept $V_{r}$ is a covariate.
Proposition 3 established that the bias of the plug-in regression coefficient $\beta$ depends on how the LLM’s error $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ covaries with $W_{r}$. This suggests a natural solution: on a random subset of the researcher’s dataset, collect $V_{r}$ using the existing measurement $f^{*}(\cdot)$. For example, the researcher reads and labels a subset of the text pieces themselves. We call text pieces on which the researcher observes $(r,W_{r},\widehat{m}(r;t),V_{r})$ the validation sample, and we refer to the remaining text pieces on which she observes $(r,W_{r},\widehat{m}(r;t))$ as the primary sample.
@@
In the validation sample, the researcher can estimate the bias $\widehat{\lambda}_{\Delta\mid W}$ by forming $\Delta_{r}=\widehat{m}(r;t)-V_{r}$ on the validation sample and regressing it on $W_{r}$. The validation sample can therefore be used for two purposes. First, the validation sample provides a target to optimize when selecting a model and prompt. For any choice of LLM $m$ and prompt $p$, the researcher can estimate the bias of the plug-in regression $\widehat{\lambda}_{\Delta\mid W}^{m,p}$ associated with the labels $\widehat{V}_{r}^{m,p}$; and thereby select the combination that results in the smallest bias.
@@
Second, and more importantly, the validation sample enables bias correction. Rather than reporting the plug-in coefficient $\widehat{\beta}$, the researcher reports
@@
This is exceedingly simple to implement: run two regressions – regress $\widehat{m}(r;t)$ on $W_{r}$ in the primary sample and regress $\Delta_{r}$ on $W_{r}$ in the validation sample – and subtract. Inference is equally straightforward: for example, researchers may bootstrap the primary and validation samples.
As discussed in Appendix C.2, the bias-corrected estimator has desirable theoretical properties. Consider a research context where the researcher randomly samples text pieces, allocating a fraction $\rho_{p}$ to the primary sample and a fraction $\rho_{v}$ to the validation sample. As the number of economically relevant text pieces grows large, the bias-corrected estimator $\widehat{\beta}^{debiased}$ is consistent for the target regression coefficient $\beta^{*}$. $\widehat{\beta}^{debiased}$ is asymptotically normal with limiting variance given by
@@
where $\sigma_{W}$ is the standard deviation of the linked variable $W_{r}$ across all text pieces, $\sigma_{\widehat{V}W}$ is the standard deviation of the product $\widehat{V}_{r}\times W_{r}$, and $\sigma_{\Delta W}$ is defined analogously. The precision of the bias-corrected estimator depends on the relative size of the validation sample versus the primary sample as well as the variability of the LLM’s label $\widehat{V}_{r}$ and measurement error $\Delta_{r}$ across text pieces.
A natural question arises: if we collect validation data anyway, why bother with the possibly mismeasured LLM labels on the primary sample? The validation-only estimator $\widehat{\beta}^{*}$ is also consistent and asymptotically normal: its limiting variance is given by $\sigma^{-4}_{W}\frac{1-\rho_{v}}{\rho_{v}}\sigma^{2}_{VW}$. Comparing these expressions reveals that the bias-corrected estimator is more precise when:
@@
This comparison depends on the relative standard deviation of the existing measurement $V_{r}$ and the LLM’s error $\Delta_{r}$. Equation (16) implies that the bias-corrected regression coefficient can be more precisely estimated than the validation-sample-only regression coefficient if the LLM’s labels are sufficiently accurate. Correctly incorporating imperfect LLM outputs can result in tighter standard errors than ignoring the LLM altogether. This phenomenon has been documented in recent machine learning research, such as AngelopoulosEtAl(23)-PPI and associated work, that combines validation data with the outputs of machine learning models to estimate downstream parameters.
In other words, a free-lunch is possible: use the LLM to solve the text processing problem while delivering precise estimates of the target regression coefficient. Importantly, LLM outputs are not substitutes for existing measurements; instead, they amplify a small validation sample.
Finally, validation data provides another critical benefit: it addresses concerns about specification searching that arise from LLM brittleness. As seen, alternative choices of prompts and models can yield dramatically different downstream estimates, creating a severe $p$-hacking risk. With countless possible prompts and multiple competing LLMs, researchers could search across specifications until finding desired results. Provided the researcher collects validation data and debiases whatever LLM output they collect, alternative choices of prompting strategy and LLM target the same empirical quantity: the parameter $\theta$ defined using the researcher’s existing measurement.
Monte Carlo Simulations based on Congressional Legislation
@@
The Congressional legislation data is well-suited to illustrate the value of validation data. The Congressional Bills Project trained teams of human annotators to label the description of each Congressional bill $r$ for its major policy topic area $V_{r}:=f^{*}(r)$, describing whether the bill falls into one of twenty possible policy areas. [^10] The Congressional Bills Project states that all annotators were trained for a full academic quarter before beginning this task (CAPCodebook). Given its widespread use, researchers are comfortable using these measurements in downstream analyses. Can an LLM automate this measurement procedure $f^{*}(\cdot)$?
@@
For a given policy topic $V_{r}$ (e.g., health, defense, etc.), linked covariate $W_{r}$ (e.g., whether the bill’s sponsor was a Democrat, etc.) and pair of large language model $m$ and prompting strategy $p$, we randomly sample $5,000$ bills. On this random sample of $5,000$ bills, we calculate the plug-in coefficient $\widehat{\beta}$ by regressing $\widehat{V}_{r}^{m,p}$ on the linked variable $W_{r}$. We next randomly reveal the existing label $V_{r}$ on $250$ (i.e. 5%) of our random sample of $5,000$ bills, which produces a validation sample. We then calculate the bias-corrected coefficient $\widehat{\beta}^{debiased}$. We repeat these steps for $1,000$ randomly sampled datasets. Across simulations, we calculate the average bias of the plug-in coefficient and the bias-corrected coefficient for the target regression $\beta^{*}$ associated with regressing the existing label $V_{r}$ on the linked variable $W_{r}$ on all $10,000$ bills. We repeat this exercise for each combination of bill topic $V_{r}$, linked covariate $W_{r}$, LLM $m$ (either GPT-3.5-Turbo, GPT-4o, GPT-5-mini and GPT-5-nano) and prompting strategy $p$. Appendix D.1 varies the size of the validation sample, finding similar results even when the validation sample contains as few as $125$ bills.